OpenStudy (anonymous):

for the function f(x)= xcubed-3xsquared+25x-75 a)find all the zeros of the function b) write the polynomial as a product of linear factors

OpenStudy (anonymous):

\[x^3-3x^2+25x-75=(x-3) \left(x^2+25\right)=(x-3)( x\pm 5 i)\]

OpenStudy (anonymous):

thats the answer?

OpenStudy (anonymous):

1. x = 3, 5i, -5i 2. (x-3)(x-5i)(x+5i)

OpenStudy (anonymous):

thanks